OpenStudy (jkbo):

Please help me A linear equation written in standard form is: 2x + 3y = 5 . Re-write the equation in slope-intercept form. Identify the slope.  A.y = 2/3x + 5/3 ; m = 2/3 B.y = -2/3x + 5/3 ; m = -2/3 C.y = 5/3x - 2/3 ; m = -2/3 D.y = -5/3x + 2/3 ; m = 2/3

OpenStudy (anonymous):

B.

OpenStudy (compassionate):

>A linear equation written in standard form is: 2x + 3y = 5 . Re-write the equation in slope-intercept form. Identify the slope.  >2x + 3y = 5 Slope intersect form is: y = mx + b, where m is the slope and b is the y-intercept. We basically solve for y. 2x + 3y = 5 3y = -2x + 5 (subtracting 2x) Now divide by 3 Can you do the rest?

OpenStudy (anonymous):

y=-2/3x+5/3

OpenStudy (jkbo):

thank you all

OpenStudy (anonymous):

from the form y=mx+b m=-2/3 b=5/3

OpenStudy (compassionate):

I was talking to @jkbo , @lcacot . Do you understand now, jkbo? You basically just solve for y.

OpenStudy (jkbo):

Yes I do

OpenStudy (jkbo):

Can you help me with one more Write an equation of a line with the given slope and y-intercept. m = 1, b = 4

OpenStudy (jkbo):

y = 4x + 1 y = x + 4 y = -1x + 4 y = x - 4

OpenStudy (jkbo):

@Compassionate

OpenStudy (compassionate):

m = 1, b = 4 y = mx + b y = 1x + 4

OpenStudy (compassionate):

Now remember, since multiplying 1 by x will give the same problem, and in Algebra we like to keep things simple, we can discard the 1. y = x + 4

OpenStudy (jkbo):

ok so its B @Compassionate

OpenStudy (compassionate):

Yes.